# A von Neumann algebraic approach to Quantum Theory on curved spacetime

### v3

@@ 7.3. Conclusion
With respect to the *-isomorphisms and compressions described above, the collection of pairs $\{({\mathcal{M}}(\mathcal{O}),H_{\omega_{\mathcal{O}}})):\mathcal{O}\in\mathcal{K}(M,g)\}$ is an example of the category VN-STDFM introduced in section 3, and hence describes a locally covariant theory in the sense of section 3 of [20].
Since the Hadamard states are a subset of quasi-free states, the above conclusion is obviously valid for these states. It is well known (see [50]) that Hadamard states have well-defined expectation values of the stress-energy tensor. Since the description of Hadamard states would go beyond the scope of this work we will limit ourselves to only a few remarks. The key property of these states is given in the language of distribution theory. And this description is well adapted to the Wightman axioms (also called Gårding-Wightman axioms), i.e. field operators localized at a point of space-time. In this paper, we analyze quantum fields localized in space-time regions. Therefore, continuing the strategy described in the paper [31], we will in the next section provide conditions also leading to the description of (regular) quantum fields with a well defined average value of the stress-energy tensor operator.
@@ 8. States invariant under the action of a Killing vector field $Z$
This will of course follow if the automorphism group $\pi_{\omega_{Z}}(a)\to\pi_{\omega_{Z}}(\alpha^{Z}_{t}(a))$ where $a\in{\mathcal{A}}(M)$ is strongly continuous at 0. With $\mu_{Z}$ denoting the inner product that forms the real part of the two point function of $\omega_{Z}$ (see [30, Theorem 5.2.24]), we claim that this automorphism group will indeed exhibit such strong continuity if for all $f\in\mathscr{D}(M,E)$ we have that $\mu_{Z}(G(f)-G(f)\circ\beta_{t}^{Z},G(f)-G(f)\circ\beta_{t}^{Z})\to 0$ as $t\to 0$. Right at the outset we note that the isometries $\beta_{t}^{Z}$ commute with the Green’s operator $G$. (See for example page 226 of [15].) This ensures that the action of these isometries on $\mathscr{D}(M,E)$, lift to an action on $\{G(f)\colon f\in\mathscr{D}(M,E)\}$. We shall constantly make silent use of this fact. Since ${\mathcal{A}}(M)$ is generated by the Weyl operators and since the span of vectors of the form $\pi_{\omega_{Z}}(W(G(f)))\Omega_{\omega_{Z}}$ is dense in the GNS Hilbert space, this claim will follow if we can show that for any $f,g\in\mathscr{D}(M,E)$ we have that $\|\pi_{\omega}((W(G(f)\circ\beta^{Z}_{t})-W(G(f)))W(G(g)))\Omega_{\omega}\|_{\omega}\to 0$ as $t\to 0$. Here we dropped the subscript $Z$ for the sake of simplicity. In the following we also set $\psi=G(f)$ and $\varphi=G(g)$ for the sake of simplicity. Now observe that
@@ 9. Enlarging the observable algebras
We recall, that the Orlicz space technique applied to regular fields satisfying an $H$-boundedness condition with respect to the density of the dual weight of the reference weight $\omega$ of the von Neumann algebra, leads to the space $L^{\cosh-1}({\mathfrak{M}})$. For all details see the paper [31]. The key observation to be made here, is that under such an $H$-boundedness restriction, the field operators affiliated to ${\mathfrak{M}}(M)$ which have all moments well defined may in a very natural way be embedded into this space. In other words, if the field operators $\{\phi_{\omega_{Z}(f)}\}$ satisfy an $H$-boundedness condition, then they can be canonically embedded into $L^{\cosh-1}({\mathfrak{M}}(M))$. In that case $\langle T_{ab}\rangle_{\omega_{Z}}$ exists and consequently the (semiclassical) Einstein equation would be well defined. In situations where we have a strongly continuous action $\alpha_{t}$ that not only leaves $\omega$ invariant but also corresponds to the modular automorphism group of some other fns weight, one can use results like [49, Corollary VIII.3.6] and [4, Proposition 4.2] to show that in the crossed product ${\mathfrak{M}}\rtimes_{\omega}\mathbb{R}$ this action is implemented by a unitary group with generator a product of the density of the dual weight and an operator affiliated to the centralizer of $\mathfrak{M}$}.
@@ 10. Concluding remarks
In this work, which is a continuation and an extension of our recent paper [31], we, firstly, present the AQFT framework in terms of von Neumann algebras. The presented framework is natural from a categorical point of view. Furthermore, the mentioned approach allows us to formulate the framework for AQFT without at the outset assuming the existence of a global Hilbert space. The important point to note here is that the presented approach allows the description of the entanglement phenomenon in QFT. In the next part of the paper, which is devoted to the study of differential calculus within the framework of field dynamics, we introduced the description of flows on local algebras without using the concept of “tangentially conditioned” algebras. In this way, we obtain a complementary and more simplified description of flows than that in our previous paper [31]. Here, flows for local quantum systems in curved space-time are defined directly using Killing vector fields acting on spacetime. The basic properties of quantum local flows have been shown. Furthermore, the conditions under which quantum Lie derivatives exist are provided. This was achieved by studying quantum local flows on regions of spacetime. We also showed the one may use the geometry of the underlying Lorentzian manifold itself to construct very natural quasi-free states at a local level, with the resultant net of local algebras satisfying the criteria of the category VN-STDFM, which in turn ensures that this net is an appropriate framework for a locally covariant quantum field theory. Hence the natural remaining question is whether the additivity property for quantum fields holds in the proposed description of nets of local algebras. For a deeper discussion of additivity the reader is refered to [44], [47]. We will here only note that the additivity property was very useful for the description of a quantum flows given in Theorem 4.1 in [31]. It is worth pointing out that for a quasi-free state satisfying mild regularity conditions, the additivity property will in the present setting also hold. To see this we need the concept of a partition of unity.
@@
[51, Theorem 1.11] Let $M$ be a $d$-dimensional $C^{k}$-differentiable second countable manifold and $\mathcal{W}$ any open cover. Then $M$ admits a countable partition of unity subordinate to the cover $\mathcal{W}$ with the support of each function in the t of unity compact.
@@
Let $\mathcal{O}\in\mathcal{K}(M,g)$ be given and let $\omega$ be a quasi-free state on ${\mathcal{A}}(\mathcal{O})$ of the form described in [30, Theorem 5.2.24] for which the inner product $\mu$ respects the mode of convergence introduced in [3, Definition 4.3.6]. Then there exists a countable family $\mathcal{V}_{n}$ of globally hyperbolic charts in $\mathcal{O}$ such that $\cup_{n}\mathcal{V}_{n}=\mathcal{O}$. Moreover ${\mathcal{M}}(\mathcal{O})=(\cup_{n=1}^{\infty}{\mathcal{M}}(\mathcal{V}_{n}))^{\prime\prime}$ where ${\mathcal{M}}(\mathcal{O})$ is the double commutant of $\pi_{\omega}({\mathcal{A}}(\mathcal{O}))$ and each ${\mathcal{M}}(\mathcal{V}_{n})$ is the double commutant of the copy of ${\mathcal{A}}(\mathcal{V}_{n})$ inside $\pi_{\omega}({\mathcal{A}}(\mathcal{O}))$.
@@
In the latter part of the proof we will silently make extensive use of the ideas and notation used in Section 8. A degree of familiarity with the flow of ideas in Section 8 is therefore necessary to navigate the proof. For each $p\in\mathcal{O}$ there exists a chart $\mathcal{V}_{p}\subset\mathcal{O}$ of $p$ which belongs to $\mathcal{K}(M,g)$ (see [37, Theorem 2.7]). We therefore have that $\cup_{p}\mathcal{V}_{p}=\mathcal{O}$. We know from Moretti’s notes [38, Definition 8.2] that $\mathcal{O}$ is second countable and hence that there must exist a countable subfamily $\{\mathcal{V}_{n}\colon n\in\mathbb{N}\}$ which still covers $\mathcal{O}$. For these charts it is clear that $(\cup_{n=1}^{\infty}{\mathcal{M}}(\mathcal{V}_{n}))^{\prime\prime}\subseteq{\mathcal{M}}(\mathcal{O})$. We are therefore left with the challenge of proving the converse containment. It is here that we need the concept of a partition of unity. Aside from what we noted earlier about second countability, it is also clear from Moretti’s notes (see the discussion following [38, Definition 2.23]) that globally hyperbolic Lorentzian manifolds satisfy all the prerequisites of Theorem 10.3. Hence there exists a countable partition of unity $\{x_{n}\colon n\in\mathbb{N}\}$ on $\mathcal{O}$ with compact supports subordinate to the family $\{\mathcal{V}_{n}\colon n\in\mathbb{N}\}$.
@@
(Here we used the fact that $\widetilde{G}_{\mathcal{V}_{n}}(g)=\widetilde{G}_{\mathcal{O}}(g_{ext})_{|\mathcal{V}_{n}}$.) The left and right hand side of the displayed equation above each respectively represent degenerate symplectic forms on the spaces $\mathcal{E}(\mathcal{V}_{n})=\{f\colon f\in\mathscr{D}(\mathcal{V}_{n},E)\}$ and $\mathcal{E}(\mathcal{O}:\mathcal{V}_{n})\{a\colon a\in\mathscr{D}(\mathcal{O},E),\mathrm{supp}(a)\subset\mathcal{V}_{n}\}$. By [3, Lemma 4.3.8] the map $\widetilde{G}_{\mathcal{V}_{n}}(f)\to\widetilde{G}_{\mathcal{O}}(f_{ext})$ will map $\mathrm{ker}(\widetilde{G}_{\mathcal{V}_{n}})$ into $\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ and hence induces a well-defined map from $\mathcal{E}(\mathcal{V}_{n})/\mathrm{ker}(\widetilde{G}_{\mathcal{V}_{n}})$ to the subspace $\mathcal{E}(\mathcal{O}:\mathcal{V}_{n})/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ of $\mathcal{E}(\mathcal{O})/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$. We remind the reader that the spaces $\mathcal{E}(\mathcal{V}_{n})/\mathrm{ker}(\widetilde{G}_{\mathcal{V}_{n}})$ and $\mathcal{E}(\mathcal{O})/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ are respectively canonically bijective to $\mathrm{SOL}(\mathcal{V}_{n})=\{\widetilde{G}_{\mathcal{V}_{n}}(f)\colon f\in\mathscr{D}(\mathcal{V}_{n},E)\}$ and $\mathrm{SOL}(\mathcal{O}:\mathcal{V}_{n})=\{\widetilde{G}_{\mathcal{O}}(a)\colon a\in\mathscr{D}(\mathcal{O},E),\mathrm{supp}(f)\subset\mathcal{V}_{n}\}$. In view of the equality in the previously displayed equation it is now clear that induced map from $\mathcal{E}(\mathcal{V}_{n})/\mathrm{ker}(\widetilde{G}_{\mathcal{V}_{n}})$ to the subspace $\mathcal{E}(\mathcal{O}:\mathcal{V}_{n})/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ of $\mathcal{E}(\mathcal{O})/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ is a symplectic isomorphism between these two spaces. By means of the aforementioned bijective correspondences, it may of course equivalently be regarded as a symplectic isomorphism from $\mathrm{SOL}(\mathcal{V}_{n})=\{\widetilde{G}_{\mathcal{V}_{n}}(f)\colon f\in\mathscr{D}(\mathcal{V}_{n},E)\}$ to $\mathrm{SOL}(\mathcal{O}:\mathcal{V}_{n})=\{\widetilde{G}_{\mathcal{O}}(f)\colon f\in\mathscr{D}(\mathcal{O},E)\}$. Hence the $\mathrm{CCR}$ $C^{*}$-algebras generated by these two sets are *-isomorphic. So when regarded as a subalgebra of ${\mathcal{A}}(\mathcal{O})$, we may identify ${\mathcal{A}}(\mathcal{V}_{n})$ with the $\mathrm{CCR}$ algebra generated by $\mathrm{SOL}(\mathcal{O}:\mathcal{V}_{n})$. In the appropriate representation we then take the double commutants of these algebras to get the corresponding statement for von Neumann algebras.
@@
Given any $f\in\mathscr{D}(\mathcal{O},E)$, we now set $f_{n}=f.x_{n}$. Checking shows that $f=\sum_{n=1}^{\infty}f_{n}$. Given any $f_{n}$, we know from the preceding theorem that there exists some $k\in\mathbb{N}$ with $\mathrm{supp}(f_{n})\subset\mathcal{V}_{k}$. Checking now shows that $f_{n}=(g_{n})_{ext}$ where $g_{n}=(f_{n})_{|\mathcal{V}_{k}}$. It is clear that $g_{n}$ belongs to $\mathscr{D}(\mathcal{V}_{k},E)$, and hence that the Weyl operator $W(\widetilde{G}_{\mathcal{O}}(f_{n}))$ belongs to ${\mathcal{M}}(\mathcal{V}_{k})$. It is also clear from the properties of the Weyl operators that the Weyl operator $W(\sum_{n=1}^{k}f_{n})$ is a just a multiple of the product $\Pi_{n=1}^{k}W(f_{n})$. To conclude we therefore need to show that in the representation engendered by $\omega$, $W(f)$ is the $\sigma$-strong limit of the sequence $(W(\sum_{n=1}^{k}f_{n}))_{k=1}^{\infty}$.
@@
For any compact subset $K$ of $M$, the local finiteness of the partition of unity ensures that for any $p\in K\cap\mathrm{supp}(f)$ we can find a neighbourhood $U_{p}\subset\mathcal{O}$ of $p$ with $U_{p}\cap\mathrm{supp}(f_{n})$ non-empty for at most finitely many of $f_{n}$’s. By compactness we may cover $K\cap\mathrm{supp}(f)$ with finitely many such neighbourhoods, from which it then follows that there must exist some $N_{K}\in\mathbb{N}$ such that $(\sum_{n=1}^{k}f_{n})_{|K}=f_{|K}$ for all $k\geq N_{K}$. A similar argument shows that the same is true for any of the derivatives of the $f$ and the partial sums $\sum_{n=1}^{k}f_{n}$. This of course means $(\sum_{n=1}^{k}f_{n})_{k=1}^{\infty}$ and each matching sequence of derivatives converges uniformly to $f$ each of its corresponding derivatives on any compact subset of $\mathcal{O}$. This has two consequences:

### v4

@@ 7.4. Emergent framework
With respect to the morphisms defined for the category VN-MA-SF introduced in section 3, the collection of pairs $\{({\mathcal{M}}(\mathcal{O}),\omega_{\mathcal{O}}):\mathcal{O}\in\mathcal{K}(M,g)\}$ describes a locally covariant theory in the sense of section 3 of [21].
Since the Hadamard states are a subset of quasi-free states, the above conclusion is obviously valid for these states. It is well known (see [53]) that Hadamard states have well-defined expectation values of the stress-energy tensor. Since the description of Hadamard states would go beyond the scope of this work we will limit ourselves to only a few remarks. The key property of these states is given in the language of distribution theory. And this description is well adapted to the Wightman axioms (also called Gårding-Wightman axioms), i.e. field operators localized at a point of space-time. In this paper, we analyze quantum fields localized in space-time regions. Therefore, continuing the strategy described in the paper [34], we will in the section 9 provide conditions also leading to the description of (regular) quantum fields with a well defined average value of the stress-energy tensor operator.
@@ 8. States invariant under the action of a Killing vector field $Z$
This will of course follow if the automorphism group $\pi_{\omega_{Z}}(a)\to\pi_{\omega_{Z}}(\alpha^{Z}_{t}(a))$ where $a\in{\mathcal{A}}(M)$ is strongly continuous at 0. With $\mu_{Z}$ denoting the inner product that forms the real part of the two point function of $\omega_{Z}$ (see [32, Theorem 5.2.24]), we claim that this automorphism group will indeed exhibit such strong continuity if for all $f\in\mathscr{D}(M,E)$ we have that $\mu_{Z}(G(f)-G(f)\circ\beta_{t}^{Z},G(f)-G(f)\circ\beta_{t}^{Z})\to 0$ as $t\to 0$. Right at the outset we note that the isometries $\beta_{t}^{Z}$ commute with the Green’s operator $G$. (See for example page 226 of [15].) This ensures that the action of these isometries on $\mathscr{D}(M,E)$, lift to an action on $\{G(f)\colon f\in\mathscr{D}(M,E)\}$. We shall constantly make silent use of this fact. Since ${\mathcal{A}}(M)$ is generated by the Weyl operators and since the span of vectors of the form $\pi_{\omega_{Z}}(W(G(f)))\Omega_{\omega_{Z}}$ is dense in the GNS Hilbert space, this claim will follow if we can show that for any $f,g\in\mathscr{D}(M,E)$ we have that $\|\pi_{\omega}((W(G(f)\circ\beta^{Z}_{t})-W(G(f)))W(G(g)))\Omega_{\omega}\|_{\omega}\to 0$ as $t\to 0$. Here we dropped the subscript $Z$ for the sake of simplicity. In the following we also set $\psi=G(f)$ and $\varphi=G(g)$ for the sake of simplicity. Now observe that
@@ 9. Enlarging the observable algebras
We recall, that the Orlicz space technique applied to regular fields satisfying an $H$-boundedness condition with respect to the density of the dual weight of the reference weight $\omega$ of the von Neumann algebra, leads to the space $L^{\cosh-1}({\mathfrak{M}})$. For all details see the paper [34]. The key observation to be made here, is that under such an $H$-boundedness restriction, the field operators affiliated to ${\mathfrak{M}}(M)$ which have all moments well defined may in a very natural way be embedded into this space. In other words, if the field operators $\{\phi_{\omega_{Z}(f)}\}$ satisfy an $H$-boundedness condition, then they can be canonically embedded into $L^{\cosh-1}({\mathfrak{M}}(M))$. In that case $\langle T_{ab}\rangle_{\omega_{Z}}$ exists and consequently the (semiclassical) Einstein equation would be well defined. In situations where we have a strongly continuous action $\alpha_{t}$ that not only leaves $\omega$ invariant but also corresponds to the modular automorphism group of some other fns weight, one can use results like [52, Corollary VIII.3.6] and [4, Proposition 4.2] to show that in the crossed product ${\mathfrak{M}}\rtimes_{\omega}\mathbb{R}$ this action is implemented by a unitary group with generator a product of the density of the dual weight and an operator affiliated to the centralizer of $\mathfrak{M}$}.
@@ 10. Concluding remarks
In this work, which is a continuation and an extension of our recent paper [34], we, firstly, present the AQFT framework in terms of von Neumann algebras. At this point we wish to re-emphasise that von Neumann algebras are the key component in noncommutative integration theory. The presented framework is natural from a categorical point of view. Furthermore, the mentioned approach allows us to formulate the framework for AQFT without at the outset assuming the existence of a global Hilbert space. The important point to note here is that the presented approach allows the description of the entanglement phenomenon in QFT. In the next part of the paper, which is devoted to the study of differential calculus within the framework of field dynamics, we introduced the description of flows on local algebras without using the concept of “tangentially conditioned” algebras. In this way, we obtain a complementary and more simplified description of flows than that in our previous paper [34]. Here, flows for local quantum systems in curved space-time are defined directly using Killing vector fields acting on spacetime. The basic properties of quantum local flows have been shown. Furthermore, the conditions under which quantum Lie derivatives exist are provided. This was achieved by studying quantum local flows on regions of spacetime. We also showed the one may use the geometry of the underlying Lorentzian manifold itself to construct very natural quasi-free states at a local level, with the resultant net of local algebras satisfying the criteria of the category VN-MA-SF, which in turn ensures that this net is an appropriate framework for a locally covariant quantum field theory. In summary, the result of our investigation suggests that the following is a good framework for AQFT which avoids the imposition of an a priori global Hilbert spaces and yet leaves room for a theory of entanglement at a local level:
  • Given a globally hyperbolic Lorentzian manifold $(M,G)$, to each $\mathcal{O}\in\mathcal{K}(M,g)$ there corresponds a von Neumann measure algebra $({\mathcal{M}}({\mathcal{O}}),\omega_{{\mathcal{O}}})$ in standard form (cf. Definition 3.5) where each $\omega_{{\mathcal{O}}}$ is a quasi-free state such that in terms of the morphisms as defined in Definition 3.5, the identification ${\mathcal{O}}\to({\mathcal{M}}({\mathcal{O}}),\omega_{{\mathcal{O}}})$ yields a locally covariant quantum field theory for $(M,g)$ in the sense of [9, Definition 2.1].
  • In the above framework the action of local flows of Killing vector fields is realised at the algebra level for short times in the sense described in Corollary 5.6. (In the case considered by Haag and Kastler, where $(M,g)$ is just Minkowski space, this property corresponds to the realisation of the action of Poincaré group at the algebra level.)
The natural remaining question is whether the additivity property for quantum fields holds in the proposed description of nets of local algebras. For a deeper discussion of additivity the reader is refered to [47], [50]. We will here only note that the additivity property was very useful for the description of a quantum flows given in Theorem 4.1 in [34]. It is worth pointing out that for a quasi-free state satisfying mild regularity conditions, the additivity property will in the present setting also hold. To see this we need the concept of a partition of unity.
@@
[54, Theorem 1.11] Let $M$ be a $d$-dimensional $C^{k}$-differentiable second countable manifold and $\mathcal{W}$ any open cover. Then $M$ admits a countable partition of unity subordinate to the cover $\mathcal{W}$ with the support of each function in the t of unity compact.
@@
Let $\mathcal{O}\in\mathcal{K}(M,g)$ be given and let $\omega$ be a quasi-free state on ${\mathcal{A}}(\mathcal{O})$ of the form described in [32, Theorem 5.2.24] for which the inner product $\mu$ respects the mode of convergence introduced in [3, Definition 4.3.6]. Then there exists a countable family $\mathcal{V}_{n}$ of globally hyperbolic charts in $\mathcal{O}$ such that $\cup_{n}\mathcal{V}_{n}=\mathcal{O}$. Moreover ${\mathcal{M}}(\mathcal{O})=(\cup_{n=1}^{\infty}{\mathcal{M}}(\mathcal{V}_{n}))^{\prime\prime}$ where ${\mathcal{M}}(\mathcal{O})$ is the double commutant of $\pi_{\omega}({\mathcal{A}}(\mathcal{O}))$ and each ${\mathcal{M}}(\mathcal{V}_{n})$ is the double commutant of the copy of ${\mathcal{A}}(\mathcal{V}_{n})$ inside $\pi_{\omega}({\mathcal{A}}(\mathcal{O}))$.
@@
In the latter part of the proof we will silently make extensive use of the ideas and notation used in Section 8. A degree of familiarity with the flow of ideas in Section 8 is therefore necessary to navigate the proof. For each $p\in\mathcal{O}$ there exists a chart $\mathcal{V}_{p}\subset\mathcal{O}$ of $p$ which belongs to $\mathcal{K}(M,g)$ (see [40, Theorem 2.7]). We therefore have that $\cup_{p}\mathcal{V}_{p}=\mathcal{O}$. We know from Moretti’s notes [41, Definition 8.2] that $\mathcal{O}$ is second countable and hence that there must exist a countable subfamily $\{\mathcal{V}_{n}\colon n\in\mathbb{N}\}$ which still covers $\mathcal{O}$. For these charts it is clear that $(\cup_{n=1}^{\infty}{\mathcal{M}}(\mathcal{V}_{n}))^{\prime\prime}\subseteq{\mathcal{M}}(\mathcal{O})$. We are therefore left with the challenge of proving the converse containment. It is here that we need the concept of a partition of unity. Aside from what we noted earlier about second countability, it is also clear from Moretti’s notes (see the discussion following [41, Definition 2.23]) that globally hyperbolic Lorentzian manifolds satisfy all the prerequisites of Theorem 10.3. Hence there exists a countable partition of unity $\{x_{n}\colon n\in\mathbb{N}\}$ on $\mathcal{O}$ with compact supports subordinate to the family $\{\mathcal{V}_{n}\colon n\in\mathbb{N}\}$.
@@
(Here we used the fact that $\widetilde{G}_{\mathcal{V}_{n}}(g)=\widetilde{G}_{\mathcal{O}}(g_{ext}){\upharpoonright}{\mathcal{V}_{n}}$.) The left and right hand side of the displayed equation above each respectively represent degenerate symplectic forms on the spaces $\mathcal{E}(\mathcal{V}_{n})=\{f\colon f\in\mathscr{D}(\mathcal{V}_{n},E)\}$ and $\mathcal{E}(\mathcal{O}:\mathcal{V}_{n})\{a\colon a\in\mathscr{D}(\mathcal{O},E),\mathrm{supp}(a)\subset\mathcal{V}_{n}\}$. By [3, Lemma 4.3.8] the map $\widetilde{G}_{\mathcal{V}_{n}}(f)\to\widetilde{G}_{\mathcal{O}}(f_{ext})$ will map $\mathrm{ker}(\widetilde{G}_{\mathcal{V}_{n}})$ into $\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ and hence induces a well-defined map from $\mathcal{E}(\mathcal{V}_{n})/\mathrm{ker}(\widetilde{G}_{\mathcal{V}_{n}})$ to the subspace $\mathcal{E}(\mathcal{O}:\mathcal{V}_{n})/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ of $\mathcal{E}(\mathcal{O})/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$. We remind the reader that the spaces $\mathcal{E}(\mathcal{V}_{n})/\mathrm{ker}(\widetilde{G}_{\mathcal{V}_{n}})$ and $\mathcal{E}(\mathcal{O})/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ are respectively canonically bijective to $\mathrm{SOL}(\mathcal{V}_{n})=\{\widetilde{G}_{\mathcal{V}_{n}}(f)\colon f\in\mathscr{D}(\mathcal{V}_{n},E)\}$ and $\mathrm{SOL}(\mathcal{O}:\mathcal{V}_{n})=\{\widetilde{G}_{\mathcal{O}}(a)\colon a\in\mathscr{D}(\mathcal{O},E),\mathrm{supp}(f)\subset\mathcal{V}_{n}\}$. In view of the equality in the previously displayed equation it is now clear that induced map from $\mathcal{E}(\mathcal{V}_{n})/\mathrm{ker}(\widetilde{G}_{\mathcal{V}_{n}})$ to the subspace $\mathcal{E}(\mathcal{O}:\mathcal{V}_{n})/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ of $\mathcal{E}(\mathcal{O})/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ is a symplectic isomorphism between these two spaces. By means of the aforementioned bijective correspondences, it may of course equivalently be regarded as a symplectic isomorphism from $\mathrm{SOL}(\mathcal{V}_{n})=\{\widetilde{G}_{\mathcal{V}_{n}}(f)\colon f\in\mathscr{D}(\mathcal{V}_{n},E)\}$ to $\mathrm{SOL}(\mathcal{O}:\mathcal{V}_{n})=\{\widetilde{G}_{\mathcal{O}}(f)\colon f\in\mathscr{D}(\mathcal{O},E)\}$. Hence the $\mathrm{CCR}$ $C^{*}$-algebras generated by these two sets are *-isomorphic. So when regarded as a subalgebra of ${\mathcal{A}}(\mathcal{O})$, we may identify ${\mathcal{A}}(\mathcal{V}_{n})$ with the $\mathrm{CCR}$ algebra generated by $\mathrm{SOL}(\mathcal{O}:\mathcal{V}_{n})$. In the appropriate representation we then take the double commutants of these algebras to get the corresponding statement for von Neumann algebras.
@@
Given any $f\in\mathscr{D}(\mathcal{O},E)$, we now set $f_{n}=f.x_{n}$. Checking shows that $f=\sum_{n=1}^{\infty}f_{n}$. Given any $f_{n}$, we know from the preceding theorem that there exists some $k\in\mathbb{N}$ with $\mathrm{supp}(f_{n})\subset\mathcal{V}_{k}$. Checking now shows that $f_{n}=(g_{n})_{ext}$ where $g_{n}=(f_{n}){\upharpoonright}{\mathcal{V}_{k}}$. It is clear that $g_{n}$ belongs to $\mathscr{D}(\mathcal{V}_{k},E)$, and hence that the Weyl operator $W(\widetilde{G}_{\mathcal{O}}(f_{n}))$ belongs to ${\mathcal{M}}(\mathcal{V}_{k})$. It is also clear from the properties of the Weyl operators that the Weyl operator $W(\sum_{n=1}^{k}f_{n})$ is a just a multiple of the product $\Pi_{n=1}^{k}W(f_{n})$. To conclude we therefore need to show that in the representation engendered by $\omega$, $W(f)$ is the $\sigma$-strong limit of the sequence $(W(\sum_{n=1}^{k}f_{n}))_{k=1}^{\infty}$.
@@
For any compact subset $K$ of $M$, the local finiteness of the partition of unity ensures that for any $p\in K\cap\mathrm{supp}(f)$ we can find a neighbourhood $U_{p}\subset\mathcal{O}$ of $p$ with $U_{p}\cap\mathrm{supp}(f_{n})$ non-empty for at most finitely many of $f_{n}$’s. By compactness we may cover $K\cap\mathrm{supp}(f)$ with finitely many such neighbourhoods, from which it then follows that there must exist some $N_{K}\in\mathbb{N}$ such that $(\sum_{n=1}^{k}f_{n}){\upharpoonright}{K}=f{\upharpoonright}{K}$ for all $k\geq N_{K}$. A similar argument shows that the same is true for any of the derivatives of the $f$ and the partial sums $\sum_{n=1}^{k}f_{n}$. This of course means $(\sum_{n=1}^{k}f_{n})_{k=1}^{\infty}$ and each matching sequence of derivatives converges uniformly to $f$ each of its corresponding derivatives on any compact subset of $\mathcal{O}$. This has two consequences:
@@
11. Declarations
11.1. Funding
The authors, Proff LE Labuschagne and WA Majewski, declare that they received no funding for this project.
11.2. Author contributions
Although there was some overlap, the bulk of the mathematical content was contributed by Prof LE Labuschagne and the bulk of the physical content by Prof WA Majewski.
11.3. Data Availability Statement
This manuscript has no associated data.
11.4. Conflict of Interest
The authors, Proff LE Labuschagne and WA Majewski, declare that they have no conflicts of interest regarding the subject matter or materials discussed in this manuscript.
11.5. Clinical Trial Number
Not applicable.